OpenStudy (anonymous):

How does A =13.9 degrees, cosA=16+13-1/2*4*square root of 13 A=cos^-1(16+13-1/2*4*square root of 13) A=13.9degrees Solve the triangle. This is what this problem was part of; a=1,b=4,C=60degrees I found C, which was square root of 13

OpenStudy (raden):

use the cosine rule : cosA = (b^2 + c^2 - a^2)/(2bc)

OpenStudy (raden):

yeah, u have done it.. :) just use ur calculator to get Cos^-1

OpenStudy (anonymous):

Cos^-1 is .5403023059

OpenStudy (raden):

cosA = (16+13-1)/(2*4*sqrt(13)) cosA = 0.97 A = cos^-1 (0.97) = 13.86 recheck ur calculating

OpenStudy (anonymous):

I get the 0.97 part but when I enter in A = cos^-1 (0.97), I dont get 13.86?

OpenStudy (raden):

but i got :)

OpenStudy (anonymous):

you entered in cos^-1 (0.97) in the calcualtor right?

OpenStudy (raden):

yep...

OpenStudy (anonymous):

when I entered it I get .5240932367?

OpenStudy (anonymous):

that was in Ti83, if I enter it in wolfram site I get 14.07degrees

OpenStudy (raden):

hmmm i dont understand use Ti83, i just use from calculator in my laptop. and got it